OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick another yellow marble? A. 2/15 B. 1/15 C. 1/9 D. 1/10

OpenStudy (anonymous):

Do you know how to draw a tree diagram of probabilities?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

Okay - hang on, im working on it

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Okay - it definitely cannot be c... I've got the answer but I'll take you through the correct process: 1) Tell me: What is the probability of picking up a RED marble?

OpenStudy (anonymous):

3 out of 10

OpenStudy (anonymous):

YES! well done. Now, 2) Since you have removed one marble, tell me: How many marbles in total now?!?!

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

Nope - try again - if I have 10 alltogether to start with and i remove one......... (hINT : 10 - 1!!)

OpenStudy (anonymous):

Oh i thought you meant one color. 9

OpenStudy (anonymous):

okay, sorry! So tell me what is the probability of getting a yellow: To figure this out we know that 1) The total has been reduced to 9 and 2) There are still 3 yellow marbles So .....the probability of getting a yellow is???????

OpenStudy (anonymous):

1/3 !

OpenStudy (anonymous):

3/9

OpenStudy (anonymous):

YES! as 3/9 = 1/3

OpenStudy (anonymous):

So, we know that Pr (RED) = 3/10 Pr (YELLOW) = 3/9 .............are you with me?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now, what I am about to tell you is IMPORTANT (REMEMBER THIS!!!) ** If you are picking marbles CONSECUTIVELY (ONE AFTER THE OTHER) => red AND yellow YOU MUST *****MULTIPLY**** THE PROBABILITIES TOGETHER!!! Do you promise you will remember this!?!? Calculate this now and see what you get... :)

OpenStudy (anonymous):

1/10?

OpenStudy (calculusfunctions):

Very Good Mr. Nuggets!

OpenStudy (anonymous):

CORRECT! Do you understand the process now?

OpenStudy (anonymous):

Lol! Yes i understand it now. And thank you for the help.

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

nice work @BarbaraKara

OpenStudy (anonymous):

Why thankyou! @pengembara_bumi1